OpenStudy (anonymous):

Which of the following is a solution of x2 - 7x = -5?

pooja195 (pooja195):

Do the directions specify what method to use? Quadratic Formula ? Factoring?

OpenStudy (anonymous):

quadratic formula

pooja195 (pooja195):

Ok can you put it into \[\huge~\rm~ax^2+bx+c=0 \] form?

OpenStudy (anonymous):

x^2 -7x+5=0 right?

pooja195 (pooja195):

yes now list your abc values

OpenStudy (anonymous):

a=1 b=-7 c=5

pooja195 (pooja195):

Good \[\huge~\rm~~x=\frac{ -(-7)~\pm \sqrt{(-7)^2-4(1)(5)} }{ 2(1) }\] \[\huge~\rm~~x=\frac{ 7)~\pm \sqrt{(-7)^2-4(1)(5)} }{ 2(1) }\] simplify

OpenStudy (anonymous):

a.7 plus or minus the square root of 69 divided by two b.7 plus or minus the square root of 29 divided by two. c.negative 7 plus or minus the square root of 29 divided by two d.negative 7 plus or minus the square root of 69 divided by two.

OpenStudy (anonymous):

these are the answers i have and my answer doesnt match any of them

pooja195 (pooja195):

\[\huge~\rm~(−7)^2−4(1)(5)\] solve this part only

OpenStudy (anonymous):

i got \[\frac{ 7\pm \sqrt{25} }{ 2}\]

OpenStudy (anonymous):

ok now what?

pooja195 (pooja195):

that should have been 29

OpenStudy (anonymous):

oh i got it i had to redo it and do it all together this time. so it would be \[\frac{ 7\pm \sqrt{29} }{ 2 }\]

pooja195 (pooja195):

\[\huge~\rm~ \frac{ 7\pm \sqrt{29} }{ 2}\]

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

thank you

pooja195 (pooja195):

no problem :) \(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions